Team A won the most games of any major league team. Team A won58 less than four times as many games as they lost. They played 162 regular-season games. How many wins and losses did team A have?

Answer by Boreal(15235)

L=x
win=4x-58
they add to 162
so 5x-58=162
5x=220
x=44 games lost
4x-58=176-58=118 games won